# MMEvalPro: Calibrating Multimodal Benchmarks Towards Trustworthy and Efficient Evaluation

### v1

@@ 1 Introduction
Ever since the birth of standardized testing, the credibility of its conclusions has been a significant concern. The same problem goes for the evaluation of recently popular Large Multimodal Models (LMMs) such as GPT4-o [22], Gemini-1.5 [24], Qwen-VL [2] and LLaVA [16]. One classic composition of such an evaluation is the multiple-choice question (MCQ), which includes an image, a question, possible choices, and an answer. This form of evaluation has higher usability compared to other evaluation methods, such as text-based evaluation and human evaluation. Many benchmarks [9, 17, 18, 19, 28, 29, 3, 4, 11, 23] designed for multimodal foundation models include a large portion of MCQs and are widely adopted in testing from basic to most advanced multimodal models [22, 2, 16], with some models reaching or even surpassing human scores on certain benchmarks [19, 11]. However, it is important to consider whether these evaluation results truly reflect the absolute capabilities of the models, especially when we are comparing them with human beings in the pursuit of AGI. In fact, several work such as PCA-Bench [4], MMStar [6] and MathVerse [30] have pointed out that the multimodal MCQ evaluation has intrinsic bias, which provides Large Language Models (LLMs) with shortcut to hack the question. FastV [5] finds that LMM could even achieve better performance on some benchmarks with only partial visual tokens. In this paper, we mainly study three popular multimodal benchmarks MMMU [28], ScienceQA [19] and MathVista [18].
In our preliminary Seeing-or-Not Comparison experiment, we found that LLMs could achieve high scores on the tested benchmarks without processing the visual data, attributed to possible data leakage, visual information problems being not related to answering the question or simply guessing the answer. Notably, the average performance gap between best LLM and best LMM is just 14.64%, which is even smaller than the gap within LMM themselves, revealing the unreliable problem of such evaluations.
@@
We carry out experiments and analyses involving $17$ different models and human experts. The findings indicate that MMEvalPro offers a more precise reflection of the true capabilities of the tested LMMs and poses a more demanding challenge. We find MMEvalPro is more trustworthy than the base benchmarks as the best LLM trails the best LMM by $23.09\%$ in MMEvalPro whereas the gap for previous benchmarks is just $14.64\%$. Significantly, even the most advanced models, including GPT-4o and Qwen-VL-Max, lag considerably behind human performance, with a notable gap of over $30\%$ in Genuine Accuracy (only $8.03\%$ for the base benchmarks). Our investigation into the factors contributing to the consistency gap illuminates the existing disparities and supports the evaluation credibility of MMEvalPro, supplying insights for future research endeavors.
@@ 2 Probing the Credibility of Multimodal Benchmarks
The fundamental assumption of any benchmark is that models achieving higher scores possess superior capabilities compared to those with lower scores. In this section, we question the credibility of such an assumption for existing benchmarks used to evaluate LMMs. We find that the multimodal benchmarks we tested are not trustworthy enough in either relative or absolute perspectives, which is concluded from two probing experiments: Seeing-or-Not Comparison and Answer Consistency Test. The processes are illustrated in Figure 2. We test the MCQ evaluation in three multimodal benchmarks across various domains including MMMU, ScienceQA-Image and MathVista. We provide the detail of dataset statistics in section 3.1, model inference hyper-parameters and prompts in Appendix-B.
Figure 2: Examples of the probing experiments.
Figure 3: Topic distribution of MMEvalPro’s data.
@@ 2.1 Seeing-or-Not Comparison
The figure indicates that the performance gap between LMMs and LLMs is significantly narrower than anticipated. Intuitively, one might assume that LLMs, which lack the ability to process visual information, would perform considerably worse on multimodal benchmarks. In fact, if we compare the scores between the best-performing LLM and LMM, we observe that for MMMU, the best LMM’s performance is only 1.1 times that of the best LLM. This performance gap is even smaller than the variability observed within LMMs themselves. Similar results go for the other tested benchmarks. What’s more surprising is that LLMs sometimes outperform their vision-enabled counterparts (GPT4-o without vision ability outperforms the LLaVA-1.5 series according to Figure 1) and there is not an apparent performance boundary between the two kinds of models. These results suggest that those benchmarks do not accurately reflect the true multimodal understanding capabilities of the tested models. The reason for this phenomenon is three-fold:
@@
2. Data leakage: During training, LLMs may inadvertently encounter similar questions or datasets, leading to unfair advantages. The benchmarks often derive questions from textbooks [19, 28], online education resources, and research papers [18], which are also sources for training datasets [25, 8].
@@
Figure 4: Examples from different splits in the MMEvalPro dataset.
@@ 2.2 Answer Consistency Test
An example is shown in Figure 2-(b), we set a perception question and a knowledge question to the MCQ from MathVista. For human examinees, the perception question and knowledge question are easier to answer than the original one since they are set to be the prerequisites in the solution path. We found that even the most advanced LMM such as GPT4o struggles at answering all related questions even if it answers the original question correctly, which is easy for human experts. A detailed case analysis is in Figure 5.
[…]
The phenomenon raises another concern for multiple-choice question (MCQ)-based multimodal evaluation benchmarks: correctly answering a question does not necessarily indicate that the model genuinely knows how to derive the final answer. A direct solution to accurately diagnose whether the model truly has the capability to solve the question is to let humans evaluate the model’s reasoning process, as done in previous works like MathVista [18]. However, human evaluation is labor-intensive and not reproducible, which complicates the broader application of the method.
Several related works use advanced LLMs like GPT-4 to replace humans in evaluating the reasoning process of LMMs [4, 30]. These methods show a strong correlation with human judgments in corresponding tests, but they result in unstable evaluation due to updates of proprietary models and inevitable API costs. Research in using LLMs as evaluators [26, 4] also finds systematic bias and a significant gap between open-source and proprietary models in terms of evaluation agreement with human experts. These drawbacks highlight the need for an economical, easy-to-use, and calibrated evaluation method for multimodal models.
@@ 3 MMEvalPro: Calibrating Multimodal Evaluation
In this section, we delve into the detailed process of constructing the MMEvalPro benchmark dataset and elucidate our methodologies for ensuring high-quality evaluation standards. Through these efforts, the MMEvalPro framework sets a new paradigm in the assessment of multimodal models, aiming to foster both accuracy and efficiency in multimodal evaluation.
@@ 3.1 Data Source
To enhance the diversity of our benchmark data, MMEvalPro integrates content from three prominent multimodal evaluation benchmarks: MMMU, ScienceQA, and MathVista. These benchmarks span educational levels from junior high to undergraduate and cover various subjects including Science, Math, and Art. The detailed descriptions of the source datasets are in Appendix A.1. Considering annotator expertise and budget, we selected $328$ questions from MMMU (dev), $1,200$ from ScienceQA-Image, and all $540$ from MathVista (testmini), totaling $2,138$ distinct multimodal MCQs.
@@ 3.2 Annotation Pipeline
1) Data Preparation: Annotators begin by thoroughly reviewing the original question to ensure a deep understanding of the underlying concepts and the solution process.
@@ 3.3 Evaluation Metrics
We propose Genuine Accuracy (GA) as the primary metric for MMEvalPro. GA equals 1 only if the model correctly answers the original question and the corresponding perception and knowledge prerequisite questions simultaneously. The second metric is the Average Accuracy (AA), which treats all questions equally and computes the average accuracy.
@@ 4.1 Setup
We test multiple LLMs and LMMs on the original benchmarks and MMEvalPro. To streamline the evaluation process, all questions are converted into a multiple-choice format, prompting the models to directly provide the answers. This approach simplifies answer matching process and eliminates the need for external models like ChatGPT, which is commonly used in previous studies such as [18, 4, 30]. In line with prior research, we incorporate an in-context demonstration for LLMs to standardize the output format. This is important as we have observed that different LLMs tend to produce varying response formats when not provided with an example. The detailed prompt and demonstration template are listed in Appendix B. For the original benchmarks, we report the average accuracy on the MCQ questions the same as we sampled for creating MMEvalPro for fair comparison. For MMEvalPro, we report both the Genuine Accuracy and Average Accuracy.
@@ 4.2 Evaluated Models
In our evaluation, we evaluate a variety of both LLMs and LMMs. For LLMs, we implement a 1-Shot setting, where a single demonstration is utilized to guide the output format. Among the LLMs assessed are some of the most advanced open-source models, including four versions of Qwen-1.5-Chat [1] with sizes spanning from 7B to 72B, as well as the LLaMA3-Instruct [25] series (8B and 70B). We also tested API-only models, such as GPT-3.5-Turbo [20] and GPT4-o [22], which rely solely on language. On the other hand, LMMs are evaluated in a zero-shot manner due to their ability to follow instructions to produce valid answer choices. In the open-source category, we tested the LLaVA-1.5 [14] and LLaVA-Next [15] series, which include models of 7B, 13B, and 34B in size. We also evaluated two of the most cutting-edge LMMs, MiniCPM-V2.5-LLaMA3-8B [10] and InternVL-1.5-Chat-20B [7], both of which have recently been made available to the public. For proprietary models, we tested GPT-4V-1106-preview [21], the latest GPT-4o [22], and Qwen-VL-Max [2]. We provide the prompts for LLMs and the hyperparameters used for LMMs in Appendix B.
@@ 4.3 Experiment Results
As shown in Table 1, we compare the performance of different models including LLMs and LMMs on MMEvalPro and the original benchmarks. We evaluated the performance of graduate students on the benchmarks as a strong baseline. The human evaluation guideline is in Appendix E. The details of human evaluation guide are in the supplement material. We also include random guess as a weak baseline performance.
We first evaluate LLMs and LMMs separately. For LLMs, all models perform poorly under the Genuine Accuracy metric in the MMEvalPro evaluation. For example, the advanced GPT-4o model achieves $67.62\%$ accuracy on the original ScienceQA-Image benchmark but drops to $10.30\%$ on its calibrated version in MMEvalPro, a decrease of $57.32$ percentage points. Similar declines are observed in MMMU (down $37.38\%$) and MathVista (down $40\%$). These low scores more accurately reflect LLMs’ general multimodal capabilities due to their lack of visual perception abilities. Open-source LLMs also show drastic declines in Genuine Accuracy. The Qwen-1.5 series suggests that larger LLMs perform better on both benchmarks, supporting the idea that stronger LLMs are better at guessing. However, this trend does not hold for the LLaMA-3 series, indicating a need for our further investigation in the future. Generally, most LLMs score below $10\%$ in Genuine Accuracy, highlighting the difficulty of the proposed benchmark for LLMs.
On the LMMs side, we also witness a large performance gap between the original benchmarks and MMEvalPro. For proprietary models, GPT-4o and Qwen-VL-Max perform the best. If we compare the performance gap of best LLM and best LMM on the original benchmark and MMEvalPro, we could find that the performance difference in scales is more clear on MMEvalPro. For example, in original MMMU benchmark, GPT-4o with vision ($54.85\%$) is only 1.1 times its non-vision LLM version ($46.09\%$). While in MMEvalPro, the LMM ($25.08\%$) version’s performance is $2.9$ times the LLM ($8.71\%$) version of GPT-4o. A similar result also goes for the other tasks in MMEvalPro. The enlarged performance discrepancy is more intuitive given the fundamental functionality difference between LLMs and LMMs. MMEvalPro could better reflect the true abilities of examinees.
@@ 4.4 Fine-grained Analysis
To better explain models’ performance on MMEvalPro, we conducted a fine-grained analysis on the experiments’ result according to the metrics proposed in section 3.3. We select the best performing open-source and proprietary LLMs and LMMs to analyze. The results are shown in Table 2. A detailed case analysis is in Figure 5. More comparisons of different models are listed in Figure 18 from the Appendix.
@@
Examining the Perception Accuracy (PA) and Knowledge Accuracy (KA) in Table 2, we find that LMMs demonstrate a greater advantage in PA compared to KA when contrasted with LLMs. This is because PA depends directly on the models’ visual capabilities, which LLMs lack. The above conclusion explains why the tested models have larger CG compared to humans, which is a potential and promising direction for future LMMs to improve on.

### v2

@@ Abstract
MMEvalPro: Calibrating Multimodal Benchmarks Towards Trustworthy and Efficient Evaluation
Jinsheng Huang1,2,3, Liang Chen1, 2, Taian Guo1,2,3, Fu Zeng4, Yusheng Zhao1,2,3 Bohan Wu1,2,3, Ye Yuan1,2,3, Haozhe Zhao1, Zhihui Guo5, Yichi Zhang1, Jingyang Yuan1,2,3 Wei Ju1,2,3, Luchen Liu1,2,3, Tianyu Liu6, Baobao Chang1, 2, Ming Zhang1, 2, 3 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University, 3PKU-Anker LLM Lab 4Chinese Academy of Medical Sciences, 5CUHK, 6Alibaba Group hjs@stu.pku.edu.cnleo.liang.chen@outlook.com{chbb,mzhang_cs}@pku.edu.cn https://mmevalpro.github.io
@@ 1 Introduction
Figure 1: Examples of the probing experiments.
Figure 2: Topic distribution of MMEvalPro’s data.
@@
Ever since the birth of standardized testing, the credibility of its conclusions has been a significant concern. The same problem goes for the evaluation of recently popular Large Multimodal Models (LMMs) such as GPT4-o (OpenAI, 2024), Gemini-1.5 (Team et al., 2024), Qwen-VL (Bai et al., 2023b) and LLaVA (Liu et al., 2023b). One classic composition of such an evaluation is the multiple-choice question (MCQ), which includes an image, a question, possible choices, and an answer. This form of evaluation has higher usability compared to other evaluation methods, such as text-based evaluation and human evaluation (Chen et al., 2024a). Many benchmarks (Fu et al., 2023; Liu et al., 2023c; Lu et al., 2024, 2022; Yue et al., 2023; Zhang et al., 2024a; Chen et al., 2023a, 2024b; Kembhavi et al., 2016; Shen et al., 2024) designed for multimodal foundation models include a large portion of MCQs and are widely adopted in testing from basic to most advanced multimodal models (OpenAI, 2024; Bai et al., 2023b; Liu et al., 2023b), with some models reaching or even surpassing human scores on certain benchmarks (Lu et al., 2022; Kembhavi et al., 2016). However, it is important to consider whether these evaluation results truly reflect the absolute capabilities of the models, especially when we are comparing them with human beings in the pursuit of AGI. In fact, several work such as PCA-Bench (Chen et al., 2024b), MMStar (Chen et al., 2024d) and MathVerse (Zhang et al., 2024b) have pointed out that the multimodal MCQ evaluation has intrinsic bias, which provides Large Language Models (LLMs) with shortcut to hack the question. FastV (Chen et al., 2024c) finds that LMM could even achieve better performance on some benchmarks with only partial visual tokens. In this paper, we mainly study three popular multimodal benchmarks MMMU (Yue et al., 2023), ScienceQA (Lu et al., 2022) and MathVista (Lu et al., 2024).
In our preliminary Seeing-or-Not Comparison experiment, we found that LLMs could achieve high scores without processing the visual data, attributed to possible data leakage, visual information problems being not related to answering the question or simply guessing the answer. Notably, the average performance gap between best LLM and best LMM is just 14.64%, which is even smaller than the gap within LMM themselves, revealing the unreliable problem of such evaluations.
@@
We carry out experiments and analyses involving $17$ different models and human experts. The findings indicate that MMEvalPro offers a more precise reflection of the tested LMMs’ capabilities and poses a more demanding challenge. MMEvalPro is more trustworthy than base benchmarks as the best LLM trails the best LMM by $23.09\%$ whereas the gap for previous benchmarks is just $14.64\%$. Significantly, even the most advanced models, including GPT-4o and Qwen-VL-Max, lag considerably behind human performance, with a notable gap of over $30\%$ in Genuine Accuracy (only $8.03\%$ for the base benchmarks). Our investigation into the factors contributing to the consistency gap illuminates the existing disparities and supports the evaluation credibility of MMEvalPro, supplying insights for future research endeavors.
@@ 2 Probing the Credibility of Multimodal Benchmarks
The fundamental assumption of any benchmark is that models achieving higher scores possess superior capabilities. In this section, we question the credibility of such an assumption for existing multimodal benchmarks. We find that the existing benchmarks are not trustworthy enough in either relative or absolute perspectives, which is concluded from two probing experiments: Seeing-or-Not Comparison and Answer Consistency Test. The processes are illustrated in Figure 1. We test the MCQ evaluation in three multimodal benchmarks across various domains including MMMU, ScienceQA-Image and MathVista. We provide the detail of dataset statistics in section 3.1, model inference hyper-parameters and prompts in Appendix-B.
@@ 2.1 Seeing-or-Not Comparison
The figure indicates that the performance gap between LMMs and LLMs is significantly narrower than anticipated. Intuitively, one might assume that LLMs, which is unable to process visual information, would perform considerably worse on multimodal benchmarks. In fact, if we compare the scores between the best-performing LLM and LMM, we observe that for MMMU, the best LMM’s performance is only 1.1 times that of the best LLM. This performance gap is even smaller than the variability observed within LMMs. Similar results go for the other tested benchmarks. It’s more surprising that LLMs sometimes outperform their vision-enabled counterparts (GPT4-o without vision ability outperforms the LLaVA-1.5 series according to Figure 3) and there is not an apparent performance boundary between the two kinds of models. These results suggest that those benchmarks do not accurately reflect the true multimodal understanding capabilities of the tested models. The reason for this phenomenon is three-fold:
@@
2. Data leakage: During training, LLMs may inadvertently encounter similar questions or datasets, leading to unfair advantages. The existing benchmarks often derive questions from textbooks (Lu et al., 2022; Yue et al., 2023), online education resources, and research papers (Lu et al., 2024), which are also sources for training datasets (Touvron et al., 2023; Clement et al., 2019).
@@
Quantitative Analysis
Dataset Proportion
MMMU 2.97%
ScienceQA 43.08%
MathVista 5.37%
Table 1: Proportions of “Image is not needed” samples in the original datasets
We analysed the proportion of samples that can be answered correctly due to “Image is not needed” in the original datasets and the results are shown in the Table 1. As the results shown, the 43.08% proportion in ScienceQA is the highest, which is also reflected in the "Not-Seeing" experiments. From analysis, we can concluded that “Image is not needed” samples significantly affected the credibility of previous benchmarks.
@@ 2.2 Answer Consistency Test
An example is shown in Figure 1-(b), we set a perception question and a knowledge question to the MCQ from MathVista. For human examinees, the perception question and knowledge question are easier to answer than the original since they are the prerequisites for the original one in the solution path. We found that even the most advanced LMM such as GPT4o struggles at answering all related questions even if it answers the original question correctly, which is easy for human experts. A detailed case analysis is in Figure 6.
The phenomenon raises another concern for multiple-choice question (MCQ)-based multimodal evaluation benchmarks: correctly answering a question does not necessarily indicate that the model genuinely knows how to derive the final answer. A direct solution to accurately diagnose whether the model truly has the capability to solve the question is to let humans evaluate the model’s reasoning process, as done in previous works like MathVista (Lu et al., 2024). However, human evaluation is labor-intensive and not reproducible, which complicates the broader application of the method.
Several related works use advanced LLMs like GPT-4 to replace humans in evaluating the reasoning process of LMMs (Chen et al., 2024b; Zhang et al., 2024b). These methods show a strong correlation with human judgments in corresponding tests, but they result in unstable evaluation due to updates of proprietary models and inevitable API costs. Research in using LLMs as evaluators (Wang et al., 2023; Chen et al., 2024b) also finds systematic bias and a significant gap between open-source and proprietary models in terms of evaluation agreement with human experts. These drawbacks highlight the need for an economical, easy-to-use, and calibrated method for multimodal models.
@@ 3 MMEvalPro: Calibrating Multimodal Evaluation
In this section, we delve into the detailed process of constructing the MMEvalPro benchmark dataset and elucidate our methodologies for ensuring high-quality evaluation standards. Through these efforts, MMEvalPro sets a new paradigm in the assessment of multimodal models, aiming to foster both accuracy and efficiency in multimodal evaluation.
@@ 3.1 Data Source
To enhance the diversity of our benchmark data, MMEvalPro integrates content from three prominent multimodal benchmarks: MMMU, ScienceQA, and MathVista. These benchmarks span educational levels from junior high to undergraduate and cover various subjects. The details of the source datasets are in Appendix A.1. Considering annotator expertise and budget, we selected $328$ questions from MMMU(dev), $1,200$ from ScienceQA-Image, and all $540$ from MathVista, totaling $2,138$ distinct multimodal MCQs.
@@ 3.2 Annotation Pipeline
1) Data Preparation: Annotators begin by thoroughly reviewing the original question to ensure a deep understanding of concepts and solutions.
@@ 3.3 Evaluation Metrics
We propose Genuine Accuracy (GA) as the primary metric for MMEvalPro. GA equals 1 only if the model correctly answers the original question and the corresponding perception and knowledge prerequisite questions simultaneously. The second metric is the Average Accuracy (AA), which computes the average accuracy of all questions.
@@ 4.1 Setup
We test multiple LLMs and LMMs on the original benchmarks and MMEvalPro. To streamline the evaluation process, all questions are converted into a multiple-choice format, prompting the models to directly provide the answers. This approach simplifies answer matching process and eliminates the need for external models like ChatGPT, which is commonly used in previous studies such as (Lu et al., 2024; Chen et al., 2024b; Zhang et al., 2024b). In line with prior research, we incorporate an in-context demonstration for LLMs to standardize the output format. This is important as we have observed that different LLMs tend to produce varying response formats when not provided with an example. The detailed prompt and demonstration template are listed in Appendix B. For the original benchmarks, we report the average accuracy on the MCQ questions the same as we sampled for creating MMEvalPro for fair comparison. For MMEvalPro, we report both the Genuine Accuracy and Average Accuracy.
@@ 4.2 Evaluated Models
In our evaluation, we evaluate a variety of both LLMs and LMMs. For LLMs, we implement a 1-Shot setting, where a single demonstration is utilized to guide the output format. Among the LLMs assessed are some of the most advanced open-source models, including four versions of Qwen-1.5-Chat (Bai et al., 2023a) with sizes spanning from 7B to 72B, as well as the LLaMA3-Instruct (Touvron et al., 2023) series (8B and 70B). We also tested API-only models, such as GPT-3.5-Turbo (OpenAI, 2022) and GPT4-o (OpenAI, 2024), which rely solely on language. On the other hand, LMMs are evaluated in a zero-shot manner due to their ability to follow instructions to produce valid answer choices. In the open-source category, we tested the LLaVA-1.5 (Liu et al., 2023a) and LLaVA-Next (Liu et al., 2024) series, which include models of 7B, 13B, and 34B in size. We also evaluated two of the most cutting-edge LMMs, MiniCPM-V2.5-LLaMA3-8B (Hu et al., 2023) and InternVL-1.5-Chat-20B (Chen et al., 2023b), both of which have recently been made available to the public. For proprietary models, we tested GPT-4V-1106-preview (OpenAI, 2023), the latest GPT-4o (OpenAI, 2024), and Qwen-VL-Max (Bai et al., 2023b). We provide the prompts for LLMs and the hyperparameters used for LMMs in Appendix B.
@@ 4.3 Experiment Results
As shown in Table 2, we compare the performance of different LLMs and LMMs on MMEvalPro and the original benchmarks. We evaluated the performance of graduate students on the benchmarks as a strong baseline. The human evaluation guideline is shown in Appendix E. We also include random guess as a weak baseline performance.
We first evaluate LLMs and LMMs separately. For LLMs, all models perform poorly under the Genuine Accuracy metric in MMEvalPro. For example, the advanced GPT-4o achieves $67.62\%$ accuracy on the original ScienceQA-Image benchmark but drops to $10.30\%$ on its calibrated version in MMEvalPro, a decrease of $57.32$ percentage points. Similar declines are observed in MMMU (down $37.38\%$) and MathVista (down $40\%$). These low scores more accurately reflect LLMs’ general multimodal capabilities due to their lack of visual perception abilities. Open-source LLMs also show drastic declines in Genuine Accuracy. The Qwen-1.5 series suggests that larger LLMs perform better on both benchmarks, supporting the idea that stronger LLMs are better at guessing. However, this trend does not hold for the LLaMA-3 series, indicating a need for our further investigation in the future. Generally, most LLMs score below $10\%$ in Genuine Accuracy, highlighting the difficulty of the proposed benchmark for LLMs.
On the LMMs side, we also witness a large performance gap between the original benchmarks and MMEvalPro. For proprietary models, GPT-4o and Qwen-VL-Max perform the best. If we compare the performance gap of best LLM and best LMM on the original benchmark and MMEvalPro, we could find that the performance difference in scales is more clear on MMEvalPro. For example, in original MMMU benchmark, GPT-4o with vision ($54.85\%$) is only 1.1 times its non-vision LLM version ($46.09\%$). While in MMEvalPro, the LMM ($25.08\%$) version’s performance is $2.9$ times the LLM ($8.71\%$) version of GPT-4o. A similar result also goes for other tasks. The enlarged performance discrepancy is more intuitive given the fundamental functionality difference between LLMs and LMMs. MMEvalPro could better reflect the true abilities of examinees.
@@ 4.4 Fine-grained Analysis
[…]
To better explain models’ performance on MMEvalPro, we conducted a fine-grained analysis on the experiments’ result according to the metrics proposed in section 3.3. We select the best performing open-source and proprietary LLMs and LMMs to analyze. The results are shown in Table 3. A detailed case analysis is shown in Figure 6. More comparisons of different models are listed in Figure 16 from the Appendix.
@@
Examining the Perception Accuracy (PA) and Knowledge Accuracy (KA) in Table 3, we find that LMMs demonstrate a greater advantage in PA compared to KA when contrasted with LLMs. This is because PA depends directly on visual capabilities, which LLMs lack. The above conclusion explains why tested models have larger CG compared to humans, which is a potential and promising direction for future LMMs to improve on.